OpenStudy (anonymous):

how to solve x^ + 6x + 5 = 0 using factoring ?

OpenStudy (anonymous):

nowyou have to factor \[(x+1)(x+5)=0\] \[x+1=0\] \[x=-1\] \[x+5=0\] \[x=-5\]

OpenStudy (anonymous):

-5